# The inducibility of Turán graphs

### v1

@@ 4.1 Preparations
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\geq 2r$ vertices, that is, $\ell=a_{1}+\cdots+a_{r}$. Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. It follows from Theorem 3.1 that $\mathrm{OPT}(F)=\{\bm{m}\}$ and
$i(F)=\kappa_{F}\cdot\frac{(m-1)_{r-1}}{m^{\ell-1}}.$
Assume that $a_{1}\geq\cdots\geq a_{r}$. If $a_{r}=1$, then the almost-balanced condition of $F$ forces $a_{1}\leq 2$, which implies $\ell\leq 2r-1$. This contradicts our assumption $\ell\geq 2r$. Therefore, we must have $a_{r}\geq 2$. Let $b_{1}>\cdots>b_{t}$ denote the distinct elements of the multiple set $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$, occurring with multiplicities $r_{1},\ldots,r_{t}$, respectively. It follows from Lemma 4.2 (ii) that
Thus, by Theorem 4.1, it suffices to verify (S1) and (S2) with some constant $\varepsilon>0$ and all sufficiently large $n$.
Fix a sufficiently large $n$. For simplicity, we will use $o(\cdot)$ to denote lower-order terms and assume that $n=km$ for some integer $k$ (i.e., $n$ is divisible by $m$). It will be clear from the proof that this divisibility assumption has no substantive effect; it is included only to allow us to omit floor and ceiling signs.
Let $G\coloneqq G_{n,\bm{m}}$ be the $n$-vertex realization with parts $V_{1},\ldots,V_{m}$, where $|V_{i}|=k$ for all $i\in[m]$. Let $\varepsilon>0$ be the constant satisfying
$1/\varepsilon=m^{\ell+1}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!.$
@@
Suppose that $\{x,y\}\subseteq V_{\alpha}$ for some $\alpha\in[m]$. Then
@@
Every part of $F$ has size at least $2$, therefore an edge inside a part $V_{\alpha}$ cannot appear in any induced copy of $F$. Hence, for $x,y$ lying in the part $V_{\alpha}$ and for all sufficiently large $n$,
$I(F,G)-I(F,G\oplus xy)$ $=\sum_{i=1}^{t}\frac{(m-1)_{r-1}\binom{k-2}{b_{i}-2}\binom{k}{b_{i}}^{r_{i}-1}\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/r_{i}}$
$=\frac{(m-1)_{r-1}\bigl(\sum_{i=1}^{r}a_{i}(a_{i}-1)\bigr)}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\cdot\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$=\frac{(m-1)_{r-1}\bigl(\sum_{i=1}^{r}a_{i}(a_{i}-1)\bigr)}{m^{\ell-2}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\cdot\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr)>\varepsilon n^{\ell-2}.$
The first equality follows by classifying the cases according to the size of the part of $F$ that contains both $x$ and $y$. Here the sum runs over all distinct part sizes $b_{i}$ of $F$. For each fixed $i$, the factor $\binom{k-2}{b_{i}-2}$ chooses the remaining $b_{i}-2$ vertices from the part $V_{\alpha}$, the factor $\binom{k}{b_{i}}^{r_{i}-1}$ chooses vertices from the other parts of size $b_{i}$, and $\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}$ handles the remaining parts. The denominator $\mathrm{sym}(a_{1},\dots,a_{r})/r_{i}$ accounts for the symmetries among parts of the same size. This proves Claim 4.7. ∎
@@
Assume without loss of generality that $x\in V_{1}$ and $y\in V_{2}$. Since $x$ and $y$ lie in different parts, the edge $xy$ is present in $G$.
We first compute the number of induced copies of $F$ in $G$ that contain the edge $xy$, which equals
$\sum_{\{i,j\}\subseteq[t]}\frac{(m-2)_{r-2}\binom{k-1}{b_{i}-1}\binom{k-1}{b_{j}-1}\binom{k}{b_{i}}^{r_{i}-1}\binom{k}{b_{j}}^{r_{j}-1}\prod_{s\in[t]\setminus\{i,j\}}\binom{k}{b_{s}}^{r_{s}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/(r_{i}r_{j})}$
$\qquad+\sum_{i\in[t],r_{i}>1}\frac{(m-2)_{r-2}\binom{k-1}{b_{i}-1}^{2}\binom{k}{b_{i}}^{r_{i}-2}\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/(r_{i}(r_{i}-1))}$
$={}$ $\frac{(m-2)_{r-2}\left(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\right)}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr)$
$={}$ $\frac{(m-2)_{r-2}\left(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\right)}{m^{\ell-2}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot(n^{\ell-2}+o(n^{\ell-2})).$ (31)
The two expressions in the leftmost side correspond to whether $x$ and $y$ lie in parts of $F$ of the same size or of different sizes. More precisely:
  • The first expression sums over distinct $i,j\in[t]$ to count copies of $F$ with $x$ in a part of size $b_{i}$ and $y$ in a part of size $b_{j}$. For a fixed pair $\{i,j\}$, the factor $\binom{k-1}{b_{i}-1}\binom{k-1}{b_{j}-1}$ chooses the remaining vertices from the parts containing $x$ and $y$, $\binom{k}{b_{i}}^{r_{i}-1}\binom{k}{b_{j}}^{r_{j}-1}$ selects vertices from the other parts of sizes $b_{i}$ and $b_{j}$, and the product $\prod_{s\notin\{i,j\}}\binom{k}{b_{s}}^{r_{s}}$ handle the remaining parts. The denominator $\mathrm{sym}(a_{1},\dots,a_{r})/(r_{i}r_{j})$ accounts for the symmetries among parts of the same size.
  • The second expression sums over $i\in[t]$ to count copies of $F$ with both $x,y$ in two distinct parts of size $b_{i}$ (clearly, $r_{i}\geq 2$). Here $\binom{k-1}{b_{i}-1}^{2}$ chooses the remaining vertices from the two particular parts of size $b_{i}$, $\binom{k}{b_{i}}^{r_{i}-2}$ selects vertices from the other $r_{i}-2$ parts of size $b_{i}$, and the remaining factors are analogous. The denominator $\mathrm{sym}(a_{1},\dots,a_{r})/(r_{i}(r_{i}-1))$ accounts for the symmetries among parts of the same size correspondingly.
@@
Suppose that $a_{r}\geq 3$. Since every part of $F$ has size at least $3$, the non‑edge $xy$ in $G\oplus xy$ cannot belong to any induced copy of $F$. Hence for sufficiently large $n$,
$I(F,G)-I(F,G\oplus xy)=\frac{(m-2)_{r-2}\left(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\right)}{m^{\ell-2}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot\bigl(n^{\ell-2}+o(n^{\ell-2})\bigr)>\varepsilon n^{\ell-2}.$
It remains to consider the case $a_{r}=2$. Note that in this case, $b_{t}=2$. So $r_{t}$ is the multiplicity of the part size $2$. Then the induced copies of $F$ in $G\oplus xy$ that contain the non-edge $xy$ correspond bijectively to the induced copies of $K_{a_{1},\ldots,a_{r-1}}$ in $G-V_{1}-V_{2}$ (i.e., the graph obtained from $G$ by deleting all vertices of $V_{1}$ and $V_{2}$). Thus the number of these copies is
$\frac{(m-2)_{r-1}}{\mathrm{sym}(a_{1},\ldots,a_{r-1})}\prod_{i=1}^{r-1}\binom{k}{a_{i}}=\frac{2(m-2)_{r-1}\cdot r_{t}}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr).$ (32)
Combining (4.3)and (32), we obtain
$I(F,G)-I(F,G\oplus xy)=\frac{(m-2)_{r-2}\left(\big(\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\big)-2(m-r)r_{t}\right)}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}(a_{j})!}\cdot\bigl(k^{\ell-2}+o(k^{\ell-2})\bigr).$
Note that $a_{i}\geq 2$ for all $i$, so $\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}\geq\sum_{\{i,j\}\subseteq[r]}2\cdot 2=4\binom{r}{2}=4r(r-1)$. Using the bound $m<\frac{4r}{3}$ from (30), we obtain
$\sum_{\{i,j\}\subseteq[r]}a_{i}a_{j}-2r_{t}(m-r)$ $\geq 4r(r-1)-2r(m-r)=2r\bigl(3r-2-m\bigr)$
$>2r\Bigl(3r-2-\frac{4r}{3}\Bigr)=2r\cdot\frac{5r-6}{3}>0\qquad(\text{since }r\geq 2).$
Therefore, $I(F,G)-I(F,G\oplus xy)>\varepsilon n^{\ell-2}$ for all sufficiently large $n$. This proves Claim 4.8. ∎
Combining Claim 4.7 and Claim 4.8, we derive that (S1) holds. Next we verify (S2).
@@
For every set $A\subseteq[m]$ with $|A|\neq m-1$,
$I(F,G_{A^{*}})-I(F,G_{A})\geq\varepsilon n^{\ell-1},$
where $A^{*}$ is any subset of $[m]$ with $|A^{*}|=m-1$.
@@
For any subset $B\subseteq[m]$ with $|B|=q$, we proceed to compute $I(F,G_{B})-I(F,G)$.
If $q=m$ or q<r1q<r-1, no induced copy of $F$ can contain $v_{B}$ (in the former case $v_{B}$ is adjacent to all parts, contradicting that $a_{r}\geq 2$; in the latter case $v_{B}$ is adjacent to fewer than r1r-1 parts, while every vertex of $F$ is adjacent to vertices from r1r-1 other parts). Otherwise, the number of such copies is given by
$\sum_{i=1}^{t}\frac{(m-q)(q)_{r-1}\binom{k}{b_{i}-1}\binom{k}{b_{i}}^{r_{i}-1}\prod_{j\in[t]\setminus\{i\}}\binom{k}{b_{j}}^{r_{j}}}{\mathrm{sym}(a_{1},\ldots,a_{r})/r_{i}}=\frac{\ell(m-q)(q)_{r-1}}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr),$
where the summation enumerates all placements of $v_{B}$ according to the size $b_{i}$ of the part of $F$ in which it lies, and we use the identity $\sum_{i}r_{i}b_{i}=\ell$. Hence we obtain
$I(F,G_{B})-I(F,G)=\frac{\ell(m-q)(q)_{r-1}}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\bigl(k^{\ell-1}+o(k^{\ell-1})\bigr).$ (33)
Define for integers $q\in[r-1,m-1]$
$H(q)\coloneqq(m-q)(q)_{r-1}.$ (34)
For convenience, we set $H(q)=0$ for all integers $q\leq r-2$. We will show that for every integer q<m1q<m-1,
$H(q)<H(m-1).$ (35)
Since $H(m-1)=(m-1)_{r-1}$, it suffices to prove that for every $q\in[r-1,m-2]$,
$(m-1)_{r-1}-(m-q)(q)_{r-1}>0.$ (36)
We prove (36) by reverse induction on $q$. For the base case q=m2q=m-2,
$(m-1)_{r-1}-2(m-2)_{r-1}=\frac{2r-1-m}{m-r}\,(m-2)_{r-1}.$
From (30) we have $m<4r/3$ and thus $2r-1-m>2r-1-4r/3=(2r-3)/3\geq 1/3$ (since $r\geq 2$). Hence the expression is positive. Now we turn to inductive step. Assume (36) holds for some $q$ with $r\leq q\leq m-2$. We now prove it for q1q-1, Using the inductive hypothesis, we have
$(m-1)_{r-1}-(m-q+1)(q-1)_{r-1}>(m-q)(q)_{r-1}-(m-q+1)(q-1)_{r-1}$
$={}$ $(q-1)_{r-1}\,\frac{(r-1)(m+1)-rq}{\,q-r+1\,}\geq(q-1)_{r-1}\,\frac{3r-m-1}{\,q-r+1\,}$,
where the last inequality follows from the assumption $q\leq m-2$. Using again $m<4r/3$ gives 3rm1>3r4r/31=(5r3)/3>03r-m-1>3r-4r/3-1=(5r-3)/3>0 for $r\geq 2$. Therefore
$(m-1)_{r-1}-(m-q+1)(q-1)_{r-1}>0,$
which is precisely (36) for q1q-1. By induction, we obtain (35).
Combining (33), (34) and (35), we obtain for all sufficiently large $n$ and any $A,A^{*}\subseteq[m]$ with $|A^{*}|=m-1$, $|A|=p\neq m-1$,
$I(F,G_{A^{*}})-I(F,G_{A})$ $=\bigl(I(F,G_{A^{*}})-I(F,G)\bigr)-\bigl(I(F,G_{A})-I(F,G)\bigr)$
$=\bigl(H(m-1)-H(p)\bigr)\cdot\frac{k^{\ell-1}+o(k^{\ell-1})}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}\geq\frac{k^{\ell-1}+o(k^{\ell-1})}{\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}$
$=\frac{n^{\ell-1}+o(n^{\ell-1})}{m^{\ell-1}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\prod_{j=1}^{r}a_{j}!}>\varepsilon n^{\ell-1}.$
Hence Claim 4.9 holds. ∎
@@
Therefore, by Theorem 4.1, the inducibility for $F$ is perfectly stable, completing Theorem 1.5. ∎
As a corollary of Theorem 1.5, for any almost balanced complete $r$-partite graph $F$ and sufficiently large $n$, all extremal graphs for $I(F,n)$ are complete multipartite. In Section 5, we strengthen this by showing that the extremal graph for $I(F,n)$ is unique, namely, the Turán graph $T_{m}(n)$ for $m=m_{r,\ell}$.
5 Exact results

### v2

@@ 4.1 Preparations
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\geq 2r$ vertices, that is, $\ell=a_{1}+\cdots+a_{r}$. Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. Since $\ell\geq 2r$, it follows from Lemma 4.2 (ii) that (29) holds.
Assume that $a_{1}\geq\cdots\geq a_{r}$. By Fact 2.7 (ii), we have $a_{r}\geq 2$. Let $b_{1}>\cdots>b_{t}$ denote the distinct elements of the multiple set $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$, occurring with multiplicities $r_{1},\ldots,r_{t}$, respectively.
It follows from Theorem 3.1 that $\mathrm{OPT}(F)=\{\bm{m}\}$. Thus, by Theorem 4.1, it suffices to verify (S1) and (S2) for some constant $\varepsilon>0$ and all sufficiently large $n$.
Fix a sufficiently large $n$ and let $k\coloneqq\lfloor n/m\rfloor$. For simplicity, we will use $o(\cdot)$ to denote lower-order terms.
Let $G\coloneqq G_{n,\bm{m}}$ be the $n$-vertex realization with parts $V_{1},\ldots,V_{m}$, where $|V_{i}|=k$ for all $i\in[m]$. Let
$\varepsilon\coloneqq\Big(4m^{\ell+1}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})\cdot a_{1}!\cdots a_{r}!\Big)^{-1}.$
@@
Suppose that $\{x,y\}\subseteq V_{i}$ for some $i\in[m]$. Then
@@
By symmetry, we may assume that $\{x,y\}\subseteq V_{1}$. Since each part of $F$ has size at least two, there is no induced copy of $F$ in $G\oplus xy$ containing $\{x,y\}$. However, it is clear that the number of induced copies of $F$ in $G$ that contain $\{x,y\}$ and with exactly $a_{1}$ vertices in $V_{1}$ is at least
$\tbinom{|V_{1}|-2}{a_{1}-2}\cdot I(K_{a_{2},\ldots,a_{r}},K[V_{2},\ldots,V_{m}]).$
By Fact 2.2, this is
$(1+o(1))\frac{1}{(a_{1}-2)!}k^{a_{1}-2}\frac{(m-1)_{r-1}}{a_{2}!\cdots a_{r!}\cdot\mathrm{sym}(a_{2},\ldots,a_{r})}k^{a_{2}+\cdots+a_{r}}>\varepsilon n^{\ell-2}.$
It follows that $I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2}$, which proves Claim 4.10. ∎
@@
By symmetry, we may assume that $(x,y)\in V_{1}\times V_{2}$. Recall that $b_{1}>\cdots>b_{t}$ are the distinct elements of the multiple set $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$, occurring with multiplicities $r_{1},\ldots,r_{t}$, respectively.
Let $\mathcal{C}$ denote the collection of all induced copies of $F$ in $G$ that contain $\{x,y\}$. For $\{i,j\}\subseteq[t]$, let $\mathcal{C}_{i,j}$ denote the collection of all induced copies of $F$ in $G$ that contain $\{x,y\}$ with $x$ lying in a part of size $b_{i}$ and $y$ lying in a part of size $b_{j}$. For each $i\in[t]$ with $r_{i}\geq 2$, let $\mathcal{C}_{i}$ denote the collection of all induced copies of $F$ in $G$ that contain $\{x,y\}$ with both $x$ and $y$ lying in parts of size $b_{i}$.
For each $\{i,j\}\subseteq[t]$, let $F_{i,j}$ denote the (r2)(r-2)-partite subgraph of $F$ obtained by removing one part of size $b_{i}$ and one part of size $b_{j}$. Then, by Fact 2.2, we have
$|\mathcal{C}_{i,j}|$ $=\tbinom{|V_{1}|-1}{b_{i}-1}\tbinom{|V_{2}|-1}{b_{j}-1}\cdot I(F_{i,j},G[V_{3},\ldots,V_{m}])$
$=(1+o(1))\frac{k^{b_{i}-1+b_{j}-1}}{(b_{i}-1)!(b_{j}-1)!}\frac{(m-2)_{r-2}}{(b_{i}!b_{j}!)^{-1}\prod_{x\in[t]}(b_{x}!)^{r_{x}}\cdot(r_{i}r_{j})^{-1}\prod_{x\in[t]}r_{x}!}k^{\ell-b_{i}-b_{j}}$
$=(1+o(1))\frac{(m-2)_{r-2}}{\prod_{x\in[t]}(b_{x}!)^{r_{x}}\cdot\prod_{x\in[t]}r_{x}!}r_{i}b_{i}r_{j}b_{j}k^{\ell-2}.$
For each $i\in[t]$ with $r_{i}\geq 2$, let $F_{i}$ denote the (r2)(r-2)-partite subgraph of $F$ obtained by removing two parts of size $b_{i}$. Then, by Fact 2.2, we have
$|\mathcal{C}_{i}|$ $=\tbinom{|V_{1}|-1}{b_{i}-1}\tbinom{|V_{2}|-1}{b_{i}-1}\cdot I(F_{i},G[V_{3},\ldots,V_{m}])$
$=(1+o(1))\frac{k^{2b_{i}-2}}{((b_{i}-1)!)^{2}}\frac{(m-2)_{r-2}}{(b_{i}!)^{-2}\prod_{x\in[t]}(b_{x}!)^{r_{x}}\cdot(r_{i}(r_{i}-1))^{-1}\prod_{x\in[t]}r_{x}!}k^{\ell-2b_{i}}$
$=(1+o(1))\frac{(m-2)_{r-2}}{\prod_{x\in[t]}(b_{x}!)^{r_{x}}\cdot\prod_{x\in[t]}r_{x}!}r_{i}(r_{i}-1)b_{i}^{2}k^{\ell-2b_{i}}.$
Therefore,
$|\mathcal{C}|$ $=\frac{(1+o(1))(m-2)_{r-2}}{\prod_{x\in[t]}(b_{x}!)^{r_{x}}\cdot\prod_{x\in[t]}r_{x}!}k^{\ell-2}\Big(\sum_{(i,j)\in([t])_{2}}r_{i}b_{i}r_{j}b_{j}+\sum_{i\in[t],r_{i}\geq 2}r_{i}(r_{i}-1)b_{i}^{2}\Big)$
$=\frac{(1+o(1))(m-2)_{r-2}}{a_{1}!\cdots a_{r}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-2}\sum_{(i,j)\in([r])_{2}}a_{i}a_{j}.$ (30)
Let $\mathcal{C}^{\prime}$ denote the collection of all induced copies of $F$ in $G\oplus xy$ that contain $\{x,y\}$. Let $\omega$ denote the number of parts of size two in $F$. Since $a_{i}\geq 2$ for every $i\in[r]$, we have (using (29))
$\sum_{(i,j)\in([r])_{2}}a_{i}a_{j}-2(m-r)\omega>4r(r-1)-2\left(\frac{4r}{3}-r\right)r=\frac{2}{3}r(5r-6).$ (31)
@@
Suppose that $\omega=0$. Since every part of $F$ has size at least $3$ and $\{x,y\}$ in not and edge in $G\oplus xy$, there is no induced copy of $F$ in $G\oplus xy$, and hence, $|\mathcal{C}^{\prime}|=0$.
Suppose that $\omega\geq 1$ (i.e. $a_{r}=2$), then $b_{t}=2$ and $r_{t}=\omega$. Note that in this case, the size of $\mathcal{C}^{\prime}$ coincides with the value of $I(K_{a_{1},\ldots,a_{r-1}},K[V_{3},\ldots,V_{m}])$, which, by Fact 2.2, is
$\frac{(1+o(1))(m-2)_{r-1}}{a_{1}!\cdots a_{r-1}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r-1})}k^{\ell-2}=\frac{(1+o(1))(m-2)_{r-2}}{a_{1}!\cdots a_{r}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-2}2(m-r)r_{t}.$
Combining this with (4.3) and (31), we obtain
$I(F,G)-I(F,G\oplus xy)$ $=\frac{(1+o(1))(m-2)_{r-2}}{a_{1}!\cdots a_{r}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-2}\Big(\sum_{(i,j)\in([r])_{2}}a_{i}a_{j}-2(m-r)\omega\Big)$
$\geq\frac{(1+o(1))2r(5r-6)(m-2)_{r-2}}{3a_{1}!\cdots a_{r}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-2}\geq\varepsilon n^{\ell-2}.$
This completes the proof of Claim 4.11. ∎
By Claims 4.10 and 4.11, (S1) holds. So it remains to verify (S2).
@@
Suppose that $B\subseteq[m]$ is a set of size $q\in[0,m]$. Then the number of induced copies of $F$ in $G_{B}$ that contain $v_{\ast}$ is
$\frac{(1+o(1))\ell\cdot H(q)}{\prod_{i\in[r]}a_{i}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-1},$
where $H\colon[m]\to\mathbb{Z}$ is the discrete function defined in Lemma 4.9.
@@
Fix a subset $B\subseteq[m]$ with $|B|=q$. Recall that $G_{B}$ is the graph obtained from $G$ by adding a new vertex $v_{\ast}$ that is adjacent to precisely all vertices in $\bigcup_{i\in B}V_{i}$. By symmetry, we may assume that $B=\{1,\ldots,q\}$.
Let $\tilde{\mathcal{C}}_{q}$ denote the collection of all induced copies of $F$ in $G_{B}$ that contain $v_{\ast}$. Since $F$ is complete $r$-partite and $a_{r}\geq 2$, if $q\in[r-2]\cup\{m\}$, then $|\tilde{\mathcal{C}}_{q}|=0$. Thus, we may assume that $q\in[r-1,m-1]$.
Recall that $b_{1}>\cdots>b_{t}$ are the distinct elements of the multiple set $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$, occurring with multiplicities $r_{1},\ldots,r_{t}$, respectively. For each $i\in[t]$, let $F_{i}^{\prime}$ denote the (r1)(r-1)-partite subgraph of $F$ by removing a part of size $b_{i}$, and let $\tilde{\mathcal{C}}_{q}^{i}$ denote the collection of all induced copies of $F$ in $G_{B}$ that contain $v_{\ast}$ such that $v_{\ast}$ lying in a part of size $b_{i}$. Then the size $\tilde{\mathcal{C}}_{q}^{i}$ coincides with the value of $\sum_{j=q+1}^{m}\binom{|V_{j}|}{b_{i}-1}\cdot I(F_{i}^{\prime},K[V_{1},\ldots,V_{q}])$, which, by Fact 2.2, is
$(1+o(1))(m-q)\frac{k^{b_{i}-1}}{(b_{i}-1)!}\cdot\frac{(q)_{r-1}}{(b_{i}!)^{-1}\prod_{i\in[r]}a_{i}!\cdot r_{i}^{-1}\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-b_{i}}=\frac{(1+o(1))(m-q)(q)_{r-1}}{\prod_{i\in[r]}a_{i}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}r_{i}b_{i}k^{\ell-1}.$
It follows that
$|\tilde{\mathcal{C}}_{q}|$ $=\frac{(1+o(1))(m-q)(q)_{r-1}}{\prod_{i\in[r]}a_{i}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-1}\cdot\sum_{i\in[t]}r_{i}b_{i}=\frac{(1+o(1))\ell(m-q)(q)_{r-1}}{\prod_{i\in[r]}a_{i}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-1}.$
This completes the proof of Claim 4.12. ∎
Fix $A\subseteq[m]$ with $|A|\neq m-1$, and let
$\Delta\coloneqq\min\big\{I(F,G_{A^{\ast}})\colon A^{\ast}\subseteq[m]\text{ and }|A^{\ast}|=m-1\big\}-I(F,G_{A}).$
It follows from Claim 4.12 and Lemma 4.9 that
$\Delta$ $=\frac{(1+o(1))\ell}{\prod_{i\in[r]}a_{i}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-1}\big(H(m-1)-H(q)\big)\geq\frac{(1+o(1))\ell}{\prod_{i\in[r]}a_{i}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}k^{\ell-1}\geq\varepsilon n^{\ell-1}.$
@@
This proves (S2). By Theorem 4.1, the inducibility problem for $F$ is perfectly stable, proving Theorem 1.5. ∎
5 Exact results and uniqueness of the extremal construction